# Location–Scale Calibration for Generalized Posterior

### v2

@@ E.1 Sampling algorithm
We describe the augmentation and Gibbs sampler used to approximate the generalized Bayes posterior $\Pi_{n}^{\eta}$ in Section 5.1. All notation is as in the main text: in particular, the Huber loss $M_{n}(\boldsymbol{\beta})$ is defined there in terms of the whitened residuals $\tilde{r}_{ij}(\boldsymbol{\beta})$ and effective scale $s_{n}=n$.
@@
$\rho_{c}(u)=\min_{t\in\mathbb{R}}\!\left\{\frac{1}{2}(u-t)^{2}+c|t|-\frac{1}{2}c^{2}\right\}.$
Hence, up to a multiplicative constant,
@@
We then use the standard normal-exponential mixture representation of the Laplace kernel: for $\kappa>0$,
$\exp(-\kappa|t|)\propto\int_{0}^{\infty}\frac{1}{\sqrt{2\pi\omega}}\exp\!\left(-\frac{t^{2}}{2\omega}\right)\frac{\kappa^{2}}{2}\exp\!\left(-\frac{\kappa^{2}}{2}\omega\right)\mathop{}\!\mathrm{d}\omega.$
Setting $\kappa=\eta c$ introduces latent scales $\omega_{ij}>0$. Under this augmentation the generalized posterior is conjugate in each block $(\boldsymbol{\beta},\boldsymbol{t},\boldsymbol{\omega})$.
Let $\boldsymbol{L}_{i}$ be the symmetric square root of the working covariance $\boldsymbol{\Sigma}_{i}$ from Section 5.1, and define $\tilde{\boldsymbol{y}}_{i}=\boldsymbol{L}_{i}^{-1}\boldsymbol{y}_{i}$ and $\boldsymbol{t}_{i}=(t_{i1},\dots,t_{in_{i}})^{\top}$. The prior is the Gaussian prior with $n$-dependent scale used in the main text, with log-density
@@
where $\rho(\boldsymbol{\beta})=2^{-1}(\boldsymbol{\beta}-\boldsymbol{\mu})^{\top}\boldsymbol{Q}(\boldsymbol{\beta}-\boldsymbol{\mu})$, $\lambda_{n}\to\lambda\in[0,\infty)$ and $r_{n}$ is a normalizing constant.
Given $(\boldsymbol{t},\boldsymbol{\omega})$, the full conditional of $\boldsymbol{\beta}$ is multivariate normal,
$\boldsymbol{\beta}\mid\boldsymbol{t},\boldsymbol{\omega},\{\mathcal{D}_{i}\}_{i=1}^{n}\sim\mathcal{N}\bigl(\boldsymbol{m}_{\mathrm{post}},\boldsymbol{\Lambda}_{\mathrm{post}}^{-1}\bigr),$
with
@@
Conditional on $(\boldsymbol{\beta},\boldsymbol{\omega})$, the latent $t_{ij}$ are independent normals. Writing $\tilde{r}_{ij}=\tilde{r}_{ij}(\boldsymbol{\beta})$,
$t_{ij}\mid\boldsymbol{\beta},\omega_{ij},\{\mathcal{D}_{i}\}_{i=1}^{n}\sim\mathcal{N}\bigl(\mu_{t,ij},\sigma_{t,ij}^{2}\bigr),\quad\sigma_{t,ij}^{2}=\bigl(\eta+\omega_{ij}^{-1}\bigr)^{-1},\quad\mu_{t,ij}=\sigma_{t,ij}^{2}\eta\tilde{r}_{ij}.$
Finally, conditional on $t_{ij}$ the latent scales $\omega_{ij}$ follow an inverse-Gaussian distribution. With the parameterization $\mathrm{IG}(\mu,\lambda)$ used in our implementation,
$\omega_{ij}\mid t_{ij}\sim\mathrm{IG}\!\left(\mu=\frac{|t_{ij}|}{\eta c},\lambda=1\right).$
@@ E.2 Simulation settings
We fix $G=100$, $n_{i}=5$ so that $n=500$, and take $p=1$ to focus on the slope $\beta$. Covariates are generated as $x_{ij}\sim\mathcal{N}(0,1)$, and the data-generating model is
@@
with independent $b_{i}\sim\mathcal{N}(0,\tau^{2})$ and $\varepsilon_{ij}\sim\mathcal{N}(0,\sigma^{2})$, with true values $\beta=2$, $\tau^{2}=2$, $\sigma^{2}=1$.
To induce model misspecification, we contaminate the errors by replacing $\varepsilon_{ij}$ with $\varepsilon_{ij}+\xi_{ij}$ with probability $0.1$, where $\xi_{ij}\sim\mathcal{N}(0,10^{2})$ independently. The loss $M_{n}$ is the Huber objective in Section 5.1 with tuning constant $c=1$ and effective scale $s_{n}=n$. For the ridge penalty $\rho(\beta)=(\beta-\mu)^{2}/2$ we set $\mu=0$ and $\lambda=0.5$. The corresponding penalized population equation (1) has solution $\beta^{\lambda}$, which we treat as the target.
@@
For each learning rate $\eta$ we compare three procedures. First, as a frequentist benchmark, we compute the Huber $M$-estimator with ridge penalty by minimizing $M_{n}(\beta)+\lambda s_{n}\rho(\beta)$ and form Wald intervals based on the sandwich variance estimator $\hat{V}_{\mathrm{target}}=\hat{J}_{\lambda}^{-1}\hat{K}\hat{J}_{\lambda}^{-1}$ in Section 5.1. Second, we compute the generalized Bayes posterior based on $M_{n}$ and the $n$-dependent Gaussian prior corresponding to the ridge penalty, using the Gibbs sampler described in Section E.1, and form equal-tailed $95\%$ credible intervals from the posterior draws of $\beta$. Third, we apply the location–scale calibration of Section 3: from the same posterior draws we estimate the working curvature $\hat{H}_{0}$ via $s_{n}\widehat{\Sigma}_{\mathrm{post}}$, estimate the target sandwich covariance $\hat{V}_{\mathrm{target}}$ by plugging in the MAP estimator and its empirical score covariance, construct the calibration operator $\hat{\Omega}$ as in (4), and transform the posterior draws to obtain calibrated credible intervals for $\beta$.

### v3

@@ E.1 Sampling algorithm
We describe the augmentation and Gibbs sampler used to approximate the generalized Bayes posterior $\Pi_{n}^{\eta}$ in Section 5.1. All notation is as in the main text: in particular, the Huber loss $M_{n}(\boldsymbol{\beta})$ is defined there in terms of the whitened residuals $\tilde{r}_{ij}(\boldsymbol{\beta})$ and the effective scale $s_{n}=n$.
@@
$\rho_{c}(u)+\frac{1}{2}c^{2}=\min_{t\in\mathbb{R}}\!\left\{\frac{1}{2}(u-t)^{2}+c|t|\right\},$
so that $\rho_{c}(0)=0$ and, for $|u|\leq c$, $\rho_{c}(u)=u^{2}/2$, while for $|u|>c$, $\rho_{c}(u)=c|u|-c^{2}/2$. Consequently, up to a multiplicative constant that does not depend on $\boldsymbol{\beta}$,
@@
We next use the standard normal-exponential mixture representation of the Laplace kernel: for $\kappa>0$,
$\exp(-\kappa|t|)\propto\int_{0}^{\infty}(2\pi\omega)^{-1/2}\exp\!\left(-\frac{t^{2}}{2\omega}\right)\frac{\kappa^{2}}{2}\exp\!\left(-\frac{\kappa^{2}}{2}\omega\right)\mathop{}\!\mathrm{d}\omega.$
Setting $\kappa=\eta c$ introduces latent scales $\omega_{ij}>0$. Under this augmentation the generalized posterior is conditionally conjugate in each block $(\boldsymbol{\beta},\boldsymbol{t},\boldsymbol{\omega})$.
Let $\boldsymbol{L}_{i}$ be the symmetric square root of the working covariance $\boldsymbol{\Sigma}_{i}$ from Section 5.1, and define $\tilde{\boldsymbol{y}}_{i}=\boldsymbol{L}_{i}^{-1}\boldsymbol{y}_{i}$ and $\boldsymbol{t}_{i}=(t_{i1},\ldots,t_{in_{i}})^{\top}$. The prior is the Gaussian prior with $n$-dependent scale used in the main text, with log-density
@@
where $\rho(\boldsymbol{\beta})=2^{-1}(\boldsymbol{\beta}-\boldsymbol{\mu})^{\top}\boldsymbol{Q}(\boldsymbol{\beta}-\boldsymbol{\mu})$, $\lambda_{n}\to\lambda\in[0,\infty)$ and $r_{n}$ is a normalizing constant that does not depend on $\boldsymbol{\beta}$.
Given $(\boldsymbol{t},\boldsymbol{\omega})$, the likelihood contribution for $\boldsymbol{\beta}$ is Gaussian in the pseudo-response $\tilde{\boldsymbol{y}}_{i}-\boldsymbol{t}_{i}$. Combining this with the Gaussian prior, the full conditional of $\boldsymbol{\beta}$ is multivariate normal,
$\boldsymbol{\beta}\mid\boldsymbol{t},\boldsymbol{\omega},\mathcal{D}\sim N\bigl(\boldsymbol{m}_{\mathrm{post}},\boldsymbol{\Lambda}_{\mathrm{post}}^{-1}\bigr),$
with precision and mean
@@
Conditional on $(\boldsymbol{\beta},\boldsymbol{\omega})$, the latent $t_{ij}$ are independent normals. Writing $\tilde{r}_{ij}=\tilde{r}_{ij}(\boldsymbol{\beta})$ for simplicity, the joint kernel in $t_{ij}$ is
$\exp\!\left\{-\frac{\eta}{2}\bigl(\tilde{r}_{ij}-t_{ij}\bigr)^{2}-\frac{t_{ij}^{2}}{2\omega_{ij}}\right\},$
which is quadratic in $t_{ij}$. Hence
$t_{ij}\mid\boldsymbol{\beta},\omega_{ij},\{\mathcal{D}_{i}\}_{i=1}^{n}\sim N\bigl(\mu_{t,ij},\sigma_{t,ij}^{2}\bigr),\quad\sigma_{t,ij}^{2}=\bigl(\eta+\omega_{ij}^{-1}\bigr)^{-1},\quad\mu_{t,ij}=\sigma_{t,ij}^{2}\eta\tilde{r}_{ij}.$
Finally, conditional on $t_{ij}$ the latent scales $\omega_{ij}$ arise only through the Laplace mixture. For each $i,j$ the conditional density of $\omega_{ij}$ is
$\omega_{ij}\mid t_{ij}\propto\omega_{ij}^{-1/2}\exp\!\left\{-\frac{1}{2}\bigl(\eta^{2}c^{2}\omega_{ij}+t_{ij}^{2}/\omega_{ij}\bigr)\right\},\quad\omega_{ij}>0,$
that is, a generalized inverse Gaussian distribution $\mathrm{GIG}(\lambda=1/2,\chi=t_{ij}^{2},\psi=(\eta c)^{2})$. Equivalently, with the inverse-Gaussian parameterization $\mathrm{IG}(\mu,\lambda)$ with density
$f(x)=\left(\frac{\lambda}{2\pi x^{3}}\right)^{1/2}\exp\!\left\{-\frac{\lambda(x-\mu)^{2}}{2\mu^{2}x}\right\},\quad x>0,$
the reciprocal variables $u_{ij}=1/\omega_{ij}$ satisfy
$u_{ij}\mid t_{ij}\sim\mathrm{IG}\!\left(\mu=\frac{\eta c}{|t_{ij}|},\lambda=\eta^{2}c^{2}\right),$
and we obtain $\omega_{ij}$ by setting $\omega_{ij}=1/u_{ij}$. This is the parameterization used in our implementation.
@@ E.2 Simulation settings
We fix $G=100$, $n_{i}=5$ so that $n=500$, and take $p=1$ to focus on the slope $\beta$. Covariates are generated as $x_{ij}\sim N(0,1)$, and the data-generating model is
@@
with independent $b_{i}\sim N(0,\tau^{2})$ and $\varepsilon_{ij}\sim N(0,\sigma^{2})$, with true values $\beta=2$, $\tau^{2}=2$, $\sigma^{2}=1$.
To induce model misspecification, we contaminate the errors by replacing $\varepsilon_{ij}$ with $\varepsilon_{ij}+\xi_{ij}$ with probability $0.1$, where $\xi_{ij}\sim N(0,10^{2})$ independently. The loss $M_{n}$ is the Huber objective in Section 5.1 with tuning constant $c=1$ and effective scale $s_{n}=n$. For the ridge penalty $\rho(\beta)=(\beta-\mu)^{2}/2$ we set $\mu=0$ and $\lambda=0.5$. The corresponding penalized population equation (1) has solution $\beta^{\lambda}$, which we treat as the target.
@@
For each learning rate $\eta$ we compare three procedures. First, as a frequentist benchmark, we compute the Huber $M$-estimator with ridge penalty by minimizing $M_{n}(\beta)+\lambda s_{n}\rho(\beta)$ and form Wald intervals based on the sandwich variance estimator $\hat{V}_{\mathrm{target}}=\hat{J}_{\lambda}^{-1}\hat{K}\hat{J}_{\lambda}^{-1}$ in Section 5.1. Second, we compute the generalized Bayes posterior based on $M_{n}$ and the $n$-dependent Gaussian prior corresponding to the ridge penalty, using the Gibbs sampler described in Section E.1, and form equal-tailed $95\%$ credible intervals from the posterior draws of $\beta$. Third, we apply the location–scale calibration of Section 3: from the same posterior draws we estimate the working curvature $\hat{H}_{0}$ via $s_{n}\hat{\Sigma}_{\mathrm{post}}$, estimate the target sandwich covariance $\hat{V}_{\mathrm{target}}$ by plugging in the MAP estimator and its empirical score covariance, construct the calibration operator $\hat{\Omega}$ as in (4), and transform the posterior draws to obtain calibrated credible intervals for $\beta$.